# On thermal transpiration and thermomolecular pressure difference

### v2

@@ 4.1. The existence of solution to (80)
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary and suppose $\Omega$ is unsealed with respect to $\partial\Omega_{1}$. Assume that $3<\gamma\leq 1$, $0\leq a<\frac{1}{4}$ and (8) holds. Then, for any $\phi_{1}\in L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3}),f_{in}\in L^{\infty}_{\alpha,0}(\partial\Omega_{1}),$ and $\phi_{2}\in L^{\infty}_{\alpha,0}(\partial\Omega_{2})$, there exists a unique solution $h\in L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$ to (80). Furthermore. the solution $h$ satisfies following estimate
$\lVert h\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\lVert h\rVert_{L^{\infty}(\partial\Omega^{+}_{2})}$ (86)
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
$\lVert A_{ij}(f)\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\leq C\zeta^{j+i-2}\Big(\epsilon\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big)$ (98)
@@
$\quad+C\left(\frac{1}{\kappa}\right)^{2}\textbf{1}_{|v|<M}\Big(\int_{\frac{\mathcal{D}}{5M}}^{\frac{\mathcal{D}}{|v|}}\int_{\mathbb{R}^{3}}\int_{1/N}^{\infty}\int_{\mathbb{R}^{3}}e^{-\frac{C_{M}}{\kappa}t}\frac{1}{s^{3}}\nu(w)f^{2}(y,w)w^{-2}(w)dwdsdydt\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (103)
@@
$\quad+\frac{C}{\kappa}\Big(\displaystyle\int_{\Omega}\int_{\frac{\mathcal{D}}{N}}^{\infty}\int_{\mathbb{R}^{3}}\frac{1}{t^{3}}\nu(u)f^{2}(y,u)w^{-2}(u)dudtdy\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (110)
@@
$\quad+\frac{C}{\zeta^{2}(1-\zeta)^{2}}\Big(\epsilon\max_{0\leq i,j\leq n-1}\lVert h^{l+1-i-j}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\{1,\frac{1}{\kappa}\}\max_{0\leq i,j\leq n-1}\lVert h^{l+1-i-j}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
Now, we construct the solution of (80) by starting from (81). It is clear that Property 4.1 still holds when $\hat{\alpha}=0$ or $\eta=1$. We rewrite (112) as
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+1+m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})},$
@@
$\leq C_{\eta}\max\{\kappa^{-1},\kappa\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\quad+C_{\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
$\quad+C_{\zeta,\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\qquad+C_{\zeta,\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\left\{\begin{aligned} &\hat{\alpha}\nu(v)\left(h^{m_{1}+1}-h^{m_{2}+1}\right)+v\cdot\nabla_{x}\left(h^{m_{1}+1}-h^{m_{2}+1}\right)+\frac{1}{\kappa}L_{w}\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=0,&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ &\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ &\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=\eta P_{w}\left(h^{m_{1}}-h^{m_{2}}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
@@
then it follows that $\lVert h^{m_{1}}-h^{m_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}<\epsilon$. From the discussion above, $\{h^{l}\}$ is a Cauchy sequence. Therefore, the limit $\lim_{l\to\infty}h^{l}=h^{\hat{\alpha},\eta}$ exists for any $\alpha,\eta$. Moreover, $h^{\hat{\alpha},\eta}$ satisfies
$\left\{\begin{aligned} \hat{\alpha}\nu(v)h^{\hat{\alpha},\eta}+v\cdot\nabla_{x}h^{\hat{\alpha},\eta}+\frac{1}{\kappa}L_{w}(h^{\hat{\alpha},\eta})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{\hat{\alpha},\eta}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{\hat{\alpha},\eta}&=\eta P_{w}(h^{\hat{\alpha},\eta})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
Moreover, we observe that the argument used in Property 4.1 also applies to $h^{\hat{\alpha},\eta}$ for any fixed $0<\eta<1$. Then we obtain
$\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (124)
$\qquad+\frac{C_{\zeta}}{\kappa}\lVert h^{\hat{\alpha},\eta}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\hat{\alpha},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{\zeta,\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
Let $0<\hat{\alpha}_{1},\hat{\alpha}_{2}<1$. We recall that
$\left\{\begin{aligned} v\cdot\nabla_{x}\left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)+\frac{1}{\kappa}L_{w}\left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)&=-\hat{\alpha}_{1}h^{\hat{\alpha}_{1},\eta}+\hat{\alpha}_{2}h^{\hat{\alpha}_{2},\eta},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ \left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ \left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right)&=\eta P_{w}\left(h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
Regard $\hat{\alpha}_{1}h^{\hat{\alpha}_{1},\eta}+\hat{\alpha}_{2}h^{\hat{\alpha}_{2},\eta}$ as $\phi_{1}$. By repeating the same process as above, we obtain
$\lVert h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (125)
$\leq C_{\zeta}\cdot\kappa\cdot\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert\langle v\rangle^{-1}h^{\hat{\alpha}_{1},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\langle v\rangle^{-1}h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}\cdot\kappa\cdot\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
This implies that $\{h^{\hat{\alpha},\eta}\}$ is a Cauchy sequence in $L^{\infty}(\Omega\times\mathbb{R}^{3})$. After we denote $h^{\eta}=\displaystyle\lim_{\alpha\to 0}h^{\hat{\alpha},\eta}$, note that $h^{\eta_{1}}-h^{\eta_{2}}$ satisfies
@@
The last inequality follows from a similar argument showing that $\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}}$ is bounded with respect to $\hat{\alpha},\eta$. Thus $\{h^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{\infty}$ and $h\coloneq\lim_{\eta\to 1}h^{\eta}$ is a mild solution of (80). The proof of the estimate (86) is similar to the proof of (124). ∎
@@
By 4.1, we obtain
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\max\{1,\kappa^{-1}\}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&\lVert m^{-1}(m_{T(x)}-m)P(f_{k})\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\kappa^{-1}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&C\lVert m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\lVert f_{k}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\kappa^{-1}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$. (128)
@@
$\lVert\Gamma(f,f)\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\leq C\lVert f\rVert^{2}_{L^{\infty}_{a,0}(\Omega\times\mathbb{R}^{3})}.$ (129)
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B})$ (131)

### v3

@@ 4.1. The existence of solution to (80)
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary and suppose $\Omega$ is unsealed with respect to $\partial\Omega_{1}$. Assume that $3<\gamma\leq 1$, $0<\alpha<\frac{1}{4}$ and (8) holds. Then, for any $\phi_{1}\in L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3}),f_{in}\in L^{\infty}_{\alpha,0}(\partial\Omega_{1}),$ and $\phi_{2}\in L^{\infty}_{\alpha,0}(\partial\Omega_{2})$, there exists a unique solution $h\in L^{\infty}_{0,0}(\Omega\times\mathbb{R}^{3})$ to (80). Furthermore. the solution $h$ satisfies following estimate
$\lVert h\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\lVert h\rVert_{L^{\infty}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3})}\Big).$ (86)
@@
$\lVert A_{ij}(f)\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\leq C\zeta^{j+i-2}\Big(\epsilon\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\left\{1,\frac{1}{\kappa}\right\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big)$ (98)
@@
$+C\left(\frac{1}{\kappa}\right)^{2}\textbf{1}_{|v|<M}\Big(\int_{\frac{\mathcal{D}}{5M}}^{\frac{\mathcal{D}}{|v|}}\int_{\mathbb{R}^{3}}\int_{1/N}^{\infty}\int_{\mathbb{R}^{3}}e^{-\frac{C_{M}}{\kappa}t}\frac{1}{s^{3}}\nu(w)f^{2}(y,w)w^{-2}(w)dwdsdydt\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\epsilon^{3-|\gamma|}_{2}+\frac{1}{1+N}+\frac{1}{N}+\frac{1}{M}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\min\left\{1,\frac{1}{\kappa}\right\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (103)
@@
$\quad+C\min\left\{1,\frac{1}{\kappa}\right\}\Big(\displaystyle\int_{\Omega}\int_{\frac{\mathcal{D}}{N}}^{\infty}\int_{\mathbb{R}^{3}}\frac{1}{t^{3}}\nu(u)f^{2}(y,u)w^{-2}(u)dudtdy\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\min\left\{1,\frac{1}{\kappa}\right\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (110)
@@
$\quad+\frac{C}{\zeta^{2}(1-\zeta)^{2}}\Big(\epsilon\max_{0\leq i,j\leq n-1}\lVert h^{l+1-i-j}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\epsilon}\min\left\{1,\frac{1}{\kappa}\right\}\max_{0\leq i,j\leq n-1}\lVert h^{l+1-i-j}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{l-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\left\{1,\frac{1}{\kappa}\right\}\max_{0\leq m\leq 2(n-1)}\lVert h^{l+1-m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
Now, we construct the solution of (80) by starting from (81). It is clear that Property 4.1 still holds when $\varpi=0$ or $\eta=1$. We rewrite (112) as
@@
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\left\{1,\frac{1}{\kappa}\right\}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+1+m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})},$
@@
$\leq C_{\eta}\max\{1,\kappa\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\quad+C_{\varpi,\eta}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
$\quad+C_{\zeta,\varpi,\eta}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\qquad+C_{\zeta,\varpi,\eta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\left\{\begin{aligned} &\varpi\nu(v)\left(h^{m_{1}+1}-h^{m_{2}+1}\right)+v\cdot\nabla_{x}\left(h^{m_{1}+1}-h^{m_{2}+1}\right)+\frac{1}{\kappa}L_{w}\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=0,&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ &\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ &\left(h^{m_{1}+1}-h^{m_{2}+1}\right)=\eta P_{w}\left(h^{m_{1}}-h^{m_{2}}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
@@
then it follows that $\lVert h^{m_{1}}-h^{m_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}<\epsilon$. From the discussion above, $\{h^{l}\}$ is a Cauchy sequence. Therefore, the limit $\lim_{l\to\infty}h^{l}=h^{\varpi,\eta}$ exists for any $\alpha,\eta$. Moreover, $h^{\varpi,\eta}$ satisfies
$\left\{\begin{aligned} \varpi\nu(v)h^{\varpi,\eta}+v\cdot\nabla_{x}h^{\varpi,\eta}+\frac{1}{\kappa}L_{w}(h^{\varpi,\eta})&=w\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ h^{\varpi,\eta}&=wf_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ h^{\varpi,\eta}&=\eta P_{w}(h^{\varpi,\eta})+w\phi_{2},&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
Moreover, we observe that the argument used in Property 4.1 also applies to $h^{\varpi,\eta}$ for any fixed $0<\eta<1$. Then we obtain
$\lVert h^{\varpi,\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (124)
$\qquad+\frac{C_{\zeta}}{\kappa}\lVert h^{\varpi,\eta}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\varpi,\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{\zeta,\varpi,\eta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
Let $0<\varpi_{1},\varpi_{2}<1$. We recall that
$\left\{\begin{aligned} v\cdot\nabla_{x}\left(h^{\varpi_{1},\eta}-h^{\varpi_{2},\eta}\right)+\frac{1}{\kappa}L_{w}\left(h^{\varpi_{1},\eta}-h^{\varpi_{2},\eta}\right)&=-\varpi_{1}h^{\varpi_{1},\eta}+\varpi_{2}h^{\varpi_{2},\eta},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ \left(h^{\varpi_{1},\eta}-h^{\varpi_{2},\eta}\right)&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ \left(h^{\varpi_{1},\eta}-h^{\varpi_{2},\eta}\right)&=\eta P_{w}\left(h^{\varpi_{1},\eta}-h^{\varpi_{2},\eta}\right),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$.
Regard $\varpi_{1}h^{\varpi_{1},\eta}+\varpi_{2}h^{\varpi_{2},\eta}$ as $\phi_{1}$. By repeating the same process as above, we obtain
$\lVert h^{\varpi_{1},\eta}-h^{\varpi_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (125)
$\leq C_{\zeta}\cdot\kappa\cdot\max\{\varpi_{1},\varpi_{2}\}\left(\lVert\langle v\rangle^{-1}h^{\varpi_{1},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\langle v\rangle^{-1}h^{\varpi_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\varpi_{1},\eta}-f^{\varpi_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}\cdot\kappa\cdot\max\{\varpi_{1},\varpi_{2}\}\left(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\varpi_{1},\eta}-f^{\varpi_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
This implies that $\{h^{\varpi,\eta}\}$ is a Cauchy sequence in $L^{\infty}(\Omega\times\mathbb{R}^{3})$. After we denote $h^{\eta}=\displaystyle\lim_{\alpha\to 0}h^{\varpi,\eta}$, note that $h^{\eta_{1}}-h^{\eta_{2}}$ satisfies
@@
The last inequality follows from a similar argument showing that $\lVert h^{\varpi,\eta}\rVert_{L^{\infty}}$ is bounded with respect to $\varpi,\eta$. Thus $\{h^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{\infty}$ and $h\coloneq\lim_{\eta\to 1}h^{\eta}$ is a mild solution of (80). The proof of the estimate (86) is similar to the proof of (124). ∎
@@
By Property 4.1, we obtain
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&\lVert m^{-1}(m_{T(x)}-m)P(f_{k})\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\kappa^{-1}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$
$\leq C_{\zeta}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&C\lVert m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\lVert f_{k}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\kappa^{-1}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$. (128)
@@
$\lVert\Gamma(f,f)\rVert_{L^{\infty}_{\alpha,-1}(\Omega\times\mathbb{R}^{3})}\leq C\lVert f\rVert^{2}_{L^{\infty}_{a,0}(\Omega\times\mathbb{R}^{3})}.$ (129)
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}(\mathcal{A}+\mathcal{B})$ (131)